Determine the FW of the following engineering project when the MARR is 15% per year. Is the project acceptable?

Project

Investment Cost

$10,000

Expected life

5 years

Market (Salvage) Value*

-$1,000

Annual Receipts

$8,000

Annual Expenses

$4,000

* A negative market value means that there is a cost to dispose of an asset.
Answer #1

PW(15%) = 10,000 + ( 8,000 - 4000)(P/A,15%,5) - 1,000 (P/F,15%,5)

                  = - 10,000 + 4000(3.3522) - 1,000 (0.4972) = 2911.6

FW(15%) = - 10,000 (F/P,15%,5) + ( 8,000 - 4000) (F/A,l5%,5) - 1,000

                       = - 10,000 (2.0114) + 4000(6.7424) - 1,000 = 5855.6

AW(15%) = 8,000 - 4000 - [ 10,000(A/P,15%,5) - (- 1000)(A/F,15%,5)]

         = 4000 - [ 10,000(0.2983) +   1,000(0.1483)] = 868.70

Since the IRR> MARR, the project is acceptable.